# EDUMATH: Generating Standards-aligned Educational Math Word Problems

### v1

@@ Appendix A Annotator Directions and Demographics
Appendix B Model Training Details and Hyperparameters
B.1 EDUMATH 12B
@@ B.3 Finetuning Ablation
Appendix C Standard Counts in Labeled ASDIV Subset
@@
Appendix D Error Analysis
@@
Using our Gemma 3 27B IT annotator’s CoT, we can also compute the frequencies of the types of errors each model makes to highlight their relative strengths/weaknesses. We can look at the proportion of overall MaC errors based on solvability, accuracy, educational appropriateness, and standards alignment as well as the proportion of each specific error type made for each criteria individually based on the response options given to human annotators reported in Appendix A (with the exception of solvability, which only has one response option). These frequencies are calculated by prompting Gemma 3 27B IT to identify the specific criteria failure and error type for that criteria in the Gemma 3 27B IT annotator’s CoT (see Appendix I.6 for the prompt used for this experiment). We show results for this experiment in Tables 6, 7, 8, and 9.
@@
As shown in the annotator directions in Appendix A, there are two primary reasons a question would be labeled as inaccurate: 1) The solution is incorrect or 2) The solution is correct but contains a reasoning error, does not make sense, or is too complicated for a student/teacher to read, both of which would be equally harmful to student learning. As shown in Table 7, all models except GPT 4.5 make a majority of their accuracy errors from the latter error type, highlighting that they tend to arrive at the right answer most of the time, even when they make an accuracy error.
@@
As shown in Appendix A, there are four primary reasons a MWP could be labeled as educationally inappropriate, including containing inappropriate material for a school setting, being strange or unrealistic, requiring no mathematical operations to solve, or being inappropriate for another reason. Of these errors, containing inappropriate material for a school setting would provide the most direct harm to young learners. As shown in Table 8, no models make this most harmful error type, which is an improvement over the results reported in Christ et al. (2024) for older models. EDUMATH 30B and the GPT models avoid educational appropriateness errors altogether, whereas the other models tend to make mistakes in outputting MWPs that are strange or unrealistic, with the exception that Gemma 3 27B IT has an equal share of MWPs that require no mathematical operations to solve.
@@
As shown in Appendix A, there are four reasons why a MWP would not be aligned with a standard or standards: 1) It is too hard for the standard(s), 2) It does not address important parts of the standard(s), 3) It does not address the standard(s) at all, or 4) It requires the use of additional math topics beyond those listed in the standard(s), each of which would be equally harmful to student learning. Table 9 shows the majority of standards alignment errors for each model are for MWPs that are missing important parts of the prespecified standard(s). Many models also frequently make MWPs that require additional math topics to solve, while it is less frequent that they make MWPs that are too hard to solve for the given standard(s) or do not address the standard(s) at all. These results suggest that models are generally able to incorporate some important elements from prespecified standard(s) when writing MWPs even when they do not fully align the problems with those standards.
Appendix E Customizing MWPs to Student Interests
@@
While Christ et al. (2024) showed it was trivial to generate MWPs aligned with student interests with LLMs, we ran an experiment to verify this for our EDUMATH models. For this experiment, we generated 1,000 MWPs from both EDUMATH models based on a randomly selected topic from a list of 188 topics 3rd-5th grade students might be interested in. We then instructed Gemma 3 27B IT to indicate whether the MWP effectively incorporated the prespecified topic, which we used to calculate the proportion of generated MWPs that successfully included a topic (see Appendix I.7 for details on the prompts and topics used in this experiment). Based on this experiment, 95.5% of EDUMATH 12B and 97.7% of EDUMATH 30B MWPs successfully included a pre-specified topic, indicating both models are effectively able to incorporate student interests.
Appendix F Student Study Details
F.1 Student Study Full Results
@@ F.2 Student Demographics
F.3 Sample Student Worksheet
@@
Appendix G Sample Annotated Data
G.1 Example that Meets All Criteria (MaC)
@@
G.2 Solvability Negative Example
@@
G.3 Accuracy Negative Example
@@
G.4 Educational Appropriateness Negative Example
@@
G.5 Standards Alignment Negative Example
@@
Appendix H Example Problem from ASDIV Subset
Below is an example 4th grade MWP from the labeled ASDIV 3rd-5th grade subset. The question comes directly from the ASDIV dataset, while the solution is generated by Llama 3.3 70B IT using the prompt in Appendix I.2.
@@ Appendix H Example Problem from ASDIV Subset
Appendix I Prompts
I.1 Standards Annotation Prompt for ASDIV Subset
@@
I.2 Solution Generation Prompt for ASDIV Subset
@@
I.3 Solution Annotation and Revision Prompts for ASDIV Subset
Figures 10 and 11 display the prompts for annotating the appropriateness of the solutions for the ASDIV 3rd-5th grade subset generated using the prompt in Appendix I.2 and revising inappropriate solutions, respectively. Both prompts are passed to Gemma 3 27B IT and the few-shot examples are consistent across prompts and were manually constructed by our research team member with K-12 teaching experience. We use a temperature of 0 for both prompts to ensure the model selects the most probable next token during generation.
@@ I.4 MWP Generation Prompt
When customizing MWPs to student interests in Section E, the formatted string "about {topic}" is added to the prompt such that the first sentence after the 8-shot examples reads, "Write a new question about {topic} based on the grade level and math topic(s) below." In this string, "{topic}" is filled in with a random selection from the list of topics reported in Appendix I.7.
@@
I.5 Gemma 3 27B IT Annotator Prompt and Sample Output
@@
I.6 Gemma 3 27B IT Prompt for Identifying Failure Types in Annotated MWPs
@@
I.7 Gemma 3 27B IT Student Interest Annotation Prompt

### v2

@@ Appendix B Annotator Directions and Demographics
Appendix C Model Training Details and Hyperparameters
C.1 EDUMATH 12B
@@ C.3 Finetuning Ablation
Appendix D Standard Counts in Labeled ASDIV Subset
@@
Appendix E Error Analysis
@@
Using our Gemma 3 27B IT annotator’s CoT, we can also compute the frequencies of the types of errors each model makes to highlight their relative strengths/weaknesses. We can look at the proportion of overall MaC errors based on solvability, accuracy, educational appropriateness, and standards alignment as well as the proportion of each specific error type made for each criteria individually based on the response options given to human annotators reported in Appendix B (with the exception of solvability, which only has one response option). These frequencies are calculated by prompting Gemma 3 27B IT to identify the specific criteria failure and error type for that criteria in the Gemma 3 27B IT annotator’s CoT (see Appendix J.6 for the prompt used for this experiment). This experiment thus allows us to pinpoint the specific error that makes each MWP not MaC, which is a framework other related work has followed Ariyarathne et al. (2025); Christ et al. (2024); Jiao et al. (2023). We show results for this experiment in Tables 6, 7, 8, and 9.
@@
As shown in the annotator directions in Appendix B, there are two primary reasons a question would be labeled as inaccurate: 1) The solution is incorrect or 2) The solution is correct but contains a reasoning error, does not make sense, or is too complicated for a student/teacher to read, both of which would be equally harmful to student learning. As shown in Table 7, all models except GPT 4.5 make a majority of their accuracy errors from the latter error type, highlighting that they tend to arrive at the right answer most of the time, even when they make an accuracy error.
@@
As shown in Appendix B, there are four primary reasons a MWP could be labeled as educationally inappropriate, including containing inappropriate material for a school setting, being strange or unrealistic, requiring no mathematical operations to solve, or being inappropriate for another reason. Of these errors, containing inappropriate material for a school setting would provide the most direct harm to young learners. As shown in Table 8, no models make this most harmful error type, which is an improvement over the results reported in Christ et al. (2024) for older models. EDUMATH 30B and the GPT models avoid educational appropriateness errors altogether, whereas the other models tend to make mistakes in outputting MWPs that are strange or unrealistic, with the exception that Gemma 3 27B IT has an equal share of MWPs that require no mathematical operations to solve.
@@
As shown in Appendix B, there are four reasons why a MWP would not be aligned with a standard or standards: 1) It is too hard for the standard(s), 2) It does not address important parts of the standard(s), 3) It does not address the standard(s) at all, or 4) It requires the use of additional math topics beyond those listed in the standard(s), each of which would be equally harmful to student learning. Table 9 shows the majority of standards alignment errors for each model are for MWPs that are missing important parts of the prespecified standard(s). Many models also frequently make MWPs that require additional math topics to solve, while it is less frequent that they make MWPs that are too hard to solve for the given standard(s) or do not address the standard(s) at all. These results suggest that models are generally able to incorporate some important elements from prespecified standard(s) when writing MWPs even when they do not fully align the problems with those standards.
Appendix F Customizing MWPs to Student Interests
@@
While Christ et al. (2024) showed it was trivial to generate MWPs aligned with student interests with LLMs, we ran an experiment to verify this for our EDUMATH models. For this experiment, we generated 1,000 MWPs from both EDUMATH models based on a randomly selected topic from a list of 188 topics 3rd-5th grade students might be interested in. We then instructed Gemma 3 27B IT to indicate whether the MWP effectively incorporated the prespecified topic, which we used to calculate the proportion of generated MWPs that successfully included a topic (see Appendix J.7 for details on the prompts and topics used in this experiment). Based on this experiment, 95.5% of EDUMATH 12B and 97.7% of EDUMATH 30B MWPs successfully included a pre-specified topic, indicating both models are effectively able to incorporate student interests.
Appendix G Student Study Details
G.1 Student Study Full Results
@@ G.2 Student Demographics
G.3 Sample Student Worksheet
@@
Appendix H Sample Annotated Data
H.1 Example that Meets All Criteria (MaC)
@@
H.2 Solvability Negative Example
@@
H.3 Accuracy Negative Example
@@
H.4 Educational Appropriateness Negative Example
@@
H.5 Standards Alignment Negative Example
@@
Appendix I Example Problem from ASDIV Subset
Below is an example 4th grade MWP from the labeled ASDIV 3rd-5th grade subset. The question comes directly from the ASDIV dataset, while the solution is generated by Llama 3.3 70B IT using the prompt in Appendix J.2.
@@ Appendix I Example Problem from ASDIV Subset
Appendix J Prompts
J.1 Standards Annotation Prompt for ASDIV Subset
@@
J.2 Solution Generation Prompt for ASDIV Subset
@@
J.3 Solution Annotation and Revision Prompts for ASDIV Subset
Figures 11 and 12 display the prompts for annotating the appropriateness of the solutions for the ASDIV 3rd-5th grade subset generated using the prompt in Appendix J.2 and revising inappropriate solutions, respectively. Both prompts are passed to Gemma 3 27B IT and the few-shot examples are consistent across prompts and were manually constructed by our research team member with K-12 teaching experience. We use a temperature of 0 for both prompts to ensure the model selects the most probable next token during generation.
@@
When customizing MWPs to student interests in Section F, the formatted string "about {topic}" is added to the prompt such that the first sentence after the 8-shot examples reads, "Write a new question about {topic} based on the grade level and math topic(s) below." In this string, "{topic}" is filled in with a random selection from the list of topics reported in Appendix J.7.
@@
J.5 Gemma 3 27B IT Annotator Prompt and Sample Output
@@
J.6 Gemma 3 27B IT Prompt for Identifying Failure Types in Annotated MWPs
@@
J.7 Gemma 3 27B IT Student Interest Annotation Prompt